# Hug Reports: Supporting Expression of Appreciation between Users and Contributors of Open Source Software Packages

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Open source softwareccs: Human-centered computing Field studies
@@ 1. Introduction
Why is appreciation rare? We suggest that many barriers to expressing appreciation stem from the fact that where users might feel appreciation (in their development environment) and what they might feel appreciation towards (a package, its modules, or its functions) is detached from where contribution activities occur (social coding platforms like GitHub) and what its units are (individual commits or pull requests). This impedes appreciation in several ways. To start with, (1) users must incur effort to navigate to existing communication channels and to locate whom to thank. Channels for communicating with contributors, such as affordances available through social coding platforms, are detached from users’ development environments. So, establishing contact with contributors requires users to leave their development environment, and navigate to communication channels elsewhere. Further, affordances of social coding platforms that allow users to direct interactions towards authors of individual commits or pull-requests quickly become ill-suited when a user wants to thank developers of the entire package or its modules. This is because a package or module may combine several commits and pull requests, and contacting authors of each atomic contribution can be prohibitively effortful. Even though social coding platforms provide other channels through which users can notify projects as a whole, (2) these channels de-emphasize appreciation because they were primarily designed to coordinate contribution activities. For instance, the primary channel through which users interact with project contributors on GitHub is by creating “Issues”, which are intended to “track ideas, feedback, tasks, or bugs for work on GitHub”. The feature’s advertised use (and name) normatively encourages communication around areas of improvement rather than appreciation. Finally, (3) when users work with a package in the development environment, the contributors and their labor are left out of focus. Today, it is possible to discover, download, and use a software package without ever learning the contributors’ identity, much less interacting with them. The impersonal way in which packages are consumed foregrounds the technical capabilities of the software while obscuring contributors’ labor, promoting an ignorant or asocial orientation towards contributors (Widder and Nafus, 2023). Because the labor of contributors is almost forgotten from the development environment, users may fail to feel a sense of personal obligation or appreciation towards contributors.
@@ 1.1. Approach
As one solution to encourage appreciation, we arrived at a conceptual design proposal for the Hug Reports system, a unidirectional communication system that would afford users the ability to select a package currently in use and direct a thanks message towards contributors, from within their code editors. The affordance would allow users to direct thanks towards entire packages (e.g. matplotlib), their modules (e.g. pyplot), or specific functions within those modules (e.g. pyplot.scatter). Our hope was that locating a communication affordance within the code editor, and allowing users to express appreciation in terms of the abstractions they were exposed to (packages, modules, or functions), would lower the effort in communicating appreciation, and that the affordance’s material presence within the editor would remind users of the contributors behind the packages they use. By explicitly accounting for the manner in which software is ultimately used in the development environment, this approach departs from prior attempts to support appreciation in open source (Overney et al., 2020; Zhang et al., 2022b). By capturing appreciation in terms of abstractions of use rather than low-level units of work, this approach is also distinct from other appreciation systems in peer production that capture appreciation towards individual units of contribution (Matias et al., [n. d.]). Given these difference from pior work, we wanted to investigate the merits of the approach. At the same time, due to our departure from previous approaches, we anticipated that unexpected cross-cutting constraints and requirements would reveal themselves when an end-to-end system is deployed.
Therefore, in this paper, we describe a field study of the Hug Reports technology probe (Figure 1). We developed a barebones version of Hug Reports as a technology probe (Hutchinson et al., 2003) to investigate: (1) how key decisions of Hug Reports affected the expression of appreciation; (2) how appreciation was felt, expressed, received, and interpreted; and (3) the opportunities that users and contributors identified for further supporting exchanges of appreciation. The probe consists of two main components: an extension for the Visual Studio Code editor through which users can express their appreciation and a notification delivered to contributors via email. The extension renders a button on every line of code that interfaces with an imported Python or JavaScript package, through which users could log a one-bit thanks and optionally a personal note, corresponding to the package (or its specific modules, classes, or functions) invoked on that line. We deployed the extension for 3 weeks with 18 developers, following which we sent notifications to the 550 contributors whose work was thanked during the course of the deployment. 26 of these contributors participated in our subsequent research activities.
@@ 1.2. Contributions
Our study had several key takeaways, including that Hug Reports encouraged appreciation that was meaningful to users and contributors, that appreciation was interpreted both as a measure of utility and as an act of expressive communication, and that contributors’ reactions to appreciation were influenced by how much they felt they had contributed to what was thanked. In addition to this, our study revealed patterns in when users expressed appreciation. Based on these findings, we discuss the opportunity and limitation of capturing appreciation towards abstractions of use (packages, modules) rather than low-level units of work (individual commits or pull-requests), and we discuss opportunities for encouraging appreciation in software development practice.
@@
  • Through the development and deployment of the Hug Reports technology probe, we provide preliminary evidence that locating a communication affordance within the code editor, and allowing users to express appreciation in terms of the abstractions they are exposed to, can encourage exchanges of appreciation towards contributors.
@@ 2. Background and Related Work
In this section, we first describe how open source software packages are developed and distributed, with a focus on the configuration of interlinking social practices and technical systems involved. Then, by considering the interaction of social and technical factors, we describe the barriers that prevent expression of appreciation. When then discuss how our work departs from prior appreciation systems. Finally, we discuss how the meaningfulness of appreciation can be preserved even as we attempt to lower the effort of expressing appreciation.
@@ 2.1. The sociotechnical system within which modern open source packages are produced and used
A central premise of CSCW research is that social practices and technical objects can’t be fully understood in isolation; they influence each other (Bowker et al., 2014). The critical interaction between social and technical factors is captured in the concept of a sociotechnical system (Cherns, 1976) where analysis is not at the level of individual technologies but instead considers the broader coherent system of technical objects and human practices. So, to understand barriers to appreciation, we begin by demarcating the sociotechnical system we are desgining within. In this work, we focus on open source software projects that are hosted on GitHub [^1] https://github.com/, and that are made available to users as packages. Here, we briefly describe how these projects are hosted, developed, and distributed with a focus on both the social practices and the technical objects.
@@
With its collaborative and social practices, contemporary open source software development occurs primarily through social coding platforms (Dabbish et al., 2012b, a), most notably GitHub. Social coding platforms provide code hosting capabilities and social features that support collaboration (Dabbish et al., 2012b). This includes features that increase visibility into development activity (Dabbish et al., 2012a) as well as channels for communication (Dabbish et al., 2012b). On GitHub, which is the most popular social coding platform (Hsieh et al., 2023; Li et al., 2021), projects are organized as repositories; a project’s repository contains its source code and hosts the conversations surrounding the project. Permissions to make commits (applying a code patch) to the hosted version of the source code are restricted to those designated as owners or collaborators on the repository. Developers who do not have such permissions, can nonetheless author code patches and work with owners or collaborators who can apply the patches on their behalf. In this work, we use the term ‘contributor’ to mean any developer who has authored a code patch that has eventually been applied to the hosted version, and the term ‘maintainer’ to refer to the subset of contributors who have commit permissions. Any developer interested in keeping track of a project, can star the repository. In each repository, developers who are contributors or users can track ideas, feedback, tasks, and bugs by creating issues. When a developer hopes to contribute a code patch to the project, they can open a pull request, a special issue that hosts conversations concerning the contribution. Activities on GitHub leave public traces. This provides visibility into software development activities at the level of actions provided by GitHub, and the underlying Git version control system (Dabbish et al., 2012b).
@@ 2.2. Barriers to expressing appreciation in open source
Within developer communities (Mayes, [n. d.]; Hammer, 2021; Lawson, 2017) as well as academic literature (Raman et al., 2020; Miller et al., 2022; Hsieh et al., 2023), there is growing recognition of how working in open source projects can often feel demotivating and stressful. Open source practitioners have shared experiences (Hammer, 2021; Lawson, 2017) about how their interactions with users can feel like a “constant stream of negativity” (Lawson, 2017). One reason why interactions can feel overwhelmingly negative is that contributors rarely receive positive feedback and appreciation from the users who benefit from their, often voluntary, labor. Nic Crane, a maintainer of Apache Arrow, suggests: “we have lots of happy but quiet users” (Mayes, [n. d.]). With little appreciation, contributors are left facing a stream of user demands and requests, some of which are even aggressive in their tone (Raman et al., 2020; Miller et al., 2022). In many discussions, the lack of positive feedback and recognition are often discussed as a cause of demotivation and burnout (Mayes, [n. d.]; Hsieh et al., 2023; Raman et al., 2020).
@@ 2.2.1. Users must incur effort to navigate to existing communication channels and to locate whom to thank
Channels for communicating with contributors are available on GitHub, where contribution activities occur. But these are detached from users’ development environments, where the packages are used. To contact contributors, users need to leave their development environment, find the right GitHub repository, navigate to GitHub, and then either open an issue or locate the contributor’s contact details. Users are more likely to undertake this effort to report issues, since improvements can directly benefit them, than to express appreciation, from which they might not see any direct benefits. This tendency for negative feedback over positive is also observed in consumer research, showing that customers with bad experiences are more inclined to leave online reviews than those with good experiences (Han and Anderson, 2020; Anderson, 1998). Further, using GitHub’s affordances to discover and contact contributors of individual commits or pull requests is straightforward because contribution activities are organized along those low-level units. But affordances that are oriented towards commits or pull-requests quickly become ill-suited when a user wants to thank developers of the entire package or its modules. A module, for instance, may combine several commits and pull requests. So, identifying all relevant contributions, and their authors, can be prohibitively effortful. Because individual commits or pull requests are abstracted away from users when they are working with a package, we suggest it can be beneficial for appreciation systems to be oriented towards the abstractions that users are exposed to—the package or its modules.
@@
In addition to supporting communication around individual commits and pull requests, GitHub provides some support for communicating with the project as a whole through issues and stars. But the uptake of these features for communicating appreciation is limited. This is because in addition to being removed from where users are, these affordances were also designed primarily to coordinate contribution activities, with values of efficiency and productivity in mind. GitHub “Issues” [^2] https://docs.github.com/en/issues/tracking-your-work-with-issues/about-issues, for instance, are advertised as a way to “track ideas, feedback, tasks, or bugs for work on GitHub”. To support effective tracking of tasks, projects can further customize the feature to provide users default issue templates for common communication such as bug reporting or feature requests [^3] https://docs.github.com/en/communities/using-templates-to-encourage-useful-issues-and-pull-requests/configuring-issue-templates-for-your-repository. Together, the name, advertised purpose, and user defaults normatively discourage appreciation. With hesitation, motivated users still reappropriate GitHub’s features to convey appreciation, as visible in issue #2264 in the react boilerplate project [^4] https://github.com/react-boilerplate/react-boilerplate/issues/2264, which is titled: “I don’t know how to thank, and show my appreciation, to the contributors on a good way”. GitHub users have pointed out how the reaction palette available within issues and pull requests lacks emojis to convey the sentiment of “Thank you” [^5] https://github.com/orgs/community/discussions/38201. While stars can be used to convey appreciation, they are also intended to function as a bookmarking tool, and a way for the platform to learn user preferences [^6] https://docs.github.com/en/get-started/exploring-projects-on-github/saving-repositories-with-stars, which obscures the intention behind ‘starring’.
@@ 2.3. Appreciation systems in peer production
Our approach departs from prior attempts to support appreciation in open source by explicitly accounting for the manner in which software is ultimately used in the development environment. Prior work has studied several systems through which users of open source software can convey appreciation (Overney et al., 2020; Zhang et al., 2022b). This includes donation platforms such as PayPal, Patreon, and OpenCollective which projects may link to from their repositories (Overney et al., 2020). Similarly, GitHub has the ‘Sponsors’(Shimada et al., 2022) feature through which users of the platform can sponsor individual developers. The ‘Say Thanks’ [^7] https://github.com/BlitzKraft/saythanks.io project provides a link that contributors can include to the repository of a project, and that users can visit to send messages of appreciation. However, much like the GitHub features we described in the previous sections, all of these appreciation systems are detached from where software use occurs—the development environment. As a result, they present many of the barriers that we describe in the previous section. By developing an appreciation system that is fine-tuned to users’ development practices, we investigate the potential benefits of such an approach as well as cross-cutting concerns that arise in its implementation. Developing an appreciation system that connects to the users’ development environment also gives us the opportunity to investigate how appreciation is experienced and expressed by users. This allows us to further extend prior literature because prior studies tend to focus solely on experiences of the contributors receiving appreciation (Overney et al., 2020; Shimada et al., 2022)
@@
Our approach tries to enable users to express appreciation in terms of the abstractions they are exposed to—the package or its modules—rather than lower-level units of contribution such as individual commits or pull requests. This distinguishes it from some existing appreciation systems in peer production that restrict users to expressing appreciation towards individual of units of work. Consider Wikipedia’s “Thanks” system (Matias et al., [n. d.]; Goel et al., 2019), through which editors can thank each other. A ”thank” link is shown next to each edit in the history view of an article. Clicking the link triggers a notification to the author of the edit. However, having to thank individual edits can be limiting when users want to express appreciation towards a group of edits, a paragraph, or a section. One comment on the talk page for Wikipedia’s “Thanks” system notes: “Sometimes I’d like to express thanks for a group of edits — for example, when none of them is individually a big deal, but together they’re really helpful. Any chance that we could get the chance to issue a single Thanks feature notice for a group of edits?” [^8] https://en.wikipedia.org/wiki/Help_talk:Notifications/Thanks/Archive_2. Our work adapts what users can thank to the form in which they consume the artifact (packages/modules) rather than the unit of production (pull requests or individual commits). We investigate the merits of such adaptation and reveal the constraints and requirements it entails. This allows us to derive implications for designing appreciation systems in other peer production contexts where units of work may be misaligned with what users feel appreciation towards .

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Open source softwareccs: Human-centered computing Field studies
@@ 1. Introduction
Why is appreciation rare? We suggest that many barriers to expressing appreciation stem from the fact that where users might feel appreciation (in their development environment) and what they might feel appreciation towards (a package, its modules, or its functions) is detached from where contribution activities occur (social coding platforms like GitHub) and what its units are (individual commits or pull requests). This impedes appreciation in several ways. To start with, (1) users must incur effort to navigate to existing communication channels and to locate whom to thank. Channels for communicating with contributors, such as affordances available through social coding platforms, are detached from users’ development environments. So, establishing contact with contributors requires users to leave their development environment, and navigate to communication channels elsewhere. Further, affordances of social coding platforms that allow users to direct interactions towards authors of individual commits or pull requests quickly become ill-suited when a user wants to thank developers of the entire package or its modules. This is because a package or module may combine several commits and pull requests, and contacting authors of each atomic contribution can be prohibitively effortful. Even though social coding platforms provide other channels through which users can notify projects as a whole, (2) these channels de-emphasize appreciation because they were primarily designed to coordinate contribution activities. For instance, the primary channel through which users interact with project contributors on GitHub is by creating “Issues”, which are intended to “track ideas, feedback, tasks, or bugs for work on GitHub”. The feature’s advertised use (and name) normatively encourages communication around areas of improvement rather than appreciation. Finally, (3) when users work with a package in the development environment, the contributors and their labor are left out of focus. Today, it is possible to discover, download, and use a software package without ever learning the contributors’ identity, much less interacting with them. The impersonal way in which packages are consumed foregrounds the technical capabilities of the software while obscuring contributors’ labor, promoting an ignorant or asocial orientation towards contributors (Widder and Nafus, 2023). Because the labor of contributors is almost forgotten from the development environment, users may fail to feel a sense of personal obligation or appreciation towards contributors.
@@ 1.1. Approach
As one solution to encourage appreciation, we arrived at a conceptual design proposal for the Hug Reports system, a unidirectional communication system that would afford users the ability to select a package currently in use and direct a thanks message towards contributors, from within their code editors. The affordance would allow users to direct thanks towards entire packages (e.g. matplotlib), their modules (e.g. pyplot), or specific functions within those modules (e.g. pyplot.scatter). Our hope was that locating a communication affordance within the code editor and allowing users to express appreciation in terms of the abstractions they were exposed to (packages, modules, or functions), would lower the effort in communicating appreciation and that the affordance’s material presence within the editor would remind users of the contributors behind the packages they use. By explicitly accounting for the manner in which software is ultimately used in the development environment, this approach departs from prior attempts to support appreciation in open source (Overney et al., 2020; Zhang et al., 2022b). By capturing appreciation in terms of abstractions of use rather than low-level units of work, this approach is also distinct from other appreciation systems in peer production that capture appreciation towards individual units of contribution (Matias et al., [n. d.]). Given these differences from pior work, we wanted to investigate the merits of the approach. At the same time, due to our departure from previous approaches, we anticipated that unexpected cross-cutting constraints and requirements would reveal themselves when an end-to-end system is deployed.
Therefore, in this paper, we describe a field study of the Hug Reports technology probe (Figure 1). We developed a barebones version of Hug Reports as a technology probe (Hutchinson et al., 2003) to investigate: (1) how key decisions of Hug Reports affected the expression of appreciation; (2) how appreciation was felt, expressed, received, and interpreted; and (3) the opportunities that users and contributors identified for further supporting exchanges of appreciation. The probe consists of two main components: an extension for the Visual Studio Code editor through which users can express their appreciation and a notification delivered to contributors via email. The extension renders a button on every line of code that interfaces with an imported Python or JavaScript package, through which users could log a one-bit thanks and optionally a personal note, corresponding to the package (or its specific modules, classes, or functions) invoked on that line. We deployed the extension for 3 weeks with 18 developers, following which we sent notifications to the 550 contributors whose work was thanked during the deployment. 26 of these contributors participated in our subsequent research activities.
@@ 1.2. Contributions
Our study had several key takeaways, including that Hug Reports encouraged appreciation that was meaningful to users and contributors, that appreciation was interpreted both as a measure of utility and as an act of expressive communication, and that contributors’ reactions to appreciation were influenced by how much they felt they had contributed to what was thanked. In addition to this, our study revealed patterns in when users expressed appreciation. Based on these findings, we discuss the opportunity and limitation of capturing appreciation towards abstractions of use (packages, modules) rather than low-level units of work (individual commits or pull requests), and we discuss opportunities for encouraging appreciation in software development practice.
@@
  • Through the development and deployment of the Hug Reports technology probe, we provide preliminary evidence that locating a communication affordance within the code editor and allowing users to express appreciation in terms of the abstractions they are exposed to, can encourage exchanges of appreciation towards contributors.
@@ 2. Background and Related Work
In this section, we first describe how open source software packages are developed and distributed, with a focus on the configuration of interlinking social practices and technical systems involved. Then, by considering the interaction of social and technical factors, we describe the barriers that prevent expression of appreciation. We then discuss how our work departs from prior appreciation systems. Finally, we discuss how the meaningfulness of appreciation can be preserved even as we attempt to lower the effort of expressing appreciation.
@@ 2.1. The sociotechnical system within which modern open source packages are produced and used
A central premise of CSCW research is that social practices and technical objects can’t be fully understood in isolation; they influence each other (Bowker et al., 2014). The critical interaction between social and technical factors is captured in the concept of a sociotechnical system (Cherns, 1976) where analysis is not at the level of individual technologies but instead considers the broader coherent system of technical objects and human practices. So, to understand barriers to appreciation, we begin by demarcating the sociotechnical system we are designing within. In this work, we focus on open source software projects that are hosted on GitHub [^1] https://github.com/, and that are made available to users as packages. Here, we briefly describe how these projects are hosted, developed, and distributed with a focus on both the social practices and the technical objects.
@@
With its collaborative and social practices, contemporary open source software development occurs primarily through social coding platforms (Dabbish et al., 2012b, a), most notably GitHub. Social coding platforms provide code hosting capabilities and social features that support collaboration (Dabbish et al., 2012b). This includes features that increase visibility into development activity (Dabbish et al., 2012a) as well as channels for communication (Dabbish et al., 2012b). On GitHub, which is the most popular social coding platform (Hsieh et al., 2023; Li et al., 2021), projects are organized as repositories; a project’s repository contains its source code and hosts the conversations surrounding the project. Permissions to make commits (applying a code patch) to the hosted version of the source code are restricted to those designated as owners or collaborators on the repository. Developers who do not have such permissions, can nonetheless author code patches and work with owners or collaborators who can apply the patches on their behalf. In this work, we use the term ‘contributor’ to mean any developer who has authored a code patch that has eventually been applied to the hosted version, and the term ‘maintainer’ to refer to the subset of contributors who have commit permissions. Any developer interested in keeping track of a project can star the repository. In each repository, developers who are contributors or users can track ideas, feedback, tasks, and bugs by creating issues. When a developer hopes to contribute a code patch to the project, they can open a pull request, a special issue that hosts conversations concerning the contribution. Activities on GitHub leave public traces. This provides visibility into software development activities at the level of actions provided by GitHub, and the underlying Git version control system (Dabbish et al., 2012b).
@@ 2.2. Barriers to expressing appreciation in open source
Within developer communities (Mayes, [n. d.]; Hammer, 2021; Lawson, 2017) as well as academic literature (Raman et al., 2020; Miller et al., 2022; Hsieh et al., 2023), there is growing recognition of how working in open source projects can often feel demotivating and stressful. Open source practitioners have shared experiences (Hammer, 2021; Lawson, 2017) about how their interactions with users can feel like a “constant stream of negativity” (Lawson, 2017). One reason why interactions can feel overwhelmingly negative is that contributors rarely receive positive feedback and appreciation from the users who benefit from their, often voluntary, labor. Nic Crane, a maintainer of Apache Arrow, suggests: “We have lots of happy but quiet users” (Mayes, [n. d.]). With little appreciation, contributors are left facing a stream of user demands and requests, some of which are even aggressive in their tone (Raman et al., 2020; Miller et al., 2022). In many discussions, the lack of positive feedback and recognition is often discussed as a cause of demotivation and burnout (Mayes, [n. d.]; Hsieh et al., 2023; Raman et al., 2020).
@@ 2.2.1. Users must incur effort to navigate to existing communication channels and to locate whom to thank
Channels for communicating with contributors are available on GitHub, where contribution activities occur. But these are detached from users’ development environments, where the packages are used. To contact contributors, users need to leave their development environment, find the right GitHub repository, navigate to GitHub, and then either open an issue or locate the contributor’s contact details. Users are more likely to undertake this effort to report issues since improvements can directly benefit them than to express appreciation, from which they might not see any direct benefits. This tendency for negative feedback over positive is also observed in consumer research, showing that customers with bad experiences are more inclined to leave online reviews than those with good experiences (Han and Anderson, 2020; Anderson, 1998). Further, using GitHub’s affordances to discover and contact contributors of individual commits or pull requests is straightforward because contribution activities are organized along those low-level units. But affordances that are oriented towards commits or pull requests quickly become ill-suited when a user wants to thank developers of the entire package or its modules. A module, for instance, may combine several commits and pull requests. So, identifying all relevant contributions, and their authors, can be prohibitively effortful. Because individual commits or pull requests are abstracted away from users when they are working with a package, we suggest it can be beneficial for appreciation systems to be oriented towards the abstractions that users are exposed to—the package or its modules.
@@
In addition to supporting communication around individual commits and pull requests, GitHub provides some support for communicating with the project as a whole through issues and stars. However, the uptake of these features for communicating appreciation is limited. This is because, in addition to being removed from where users are, these affordances were also designed primarily to coordinate contribution activities, with values of efficiency and productivity in mind. GitHub “Issues” [^2] https://docs.github.com/en/issues/tracking-your-work-with-issues/about-issues, for instance, are advertised as a way to “track ideas, feedback, tasks, or bugs for work on GitHub”. To support effective tracking of tasks, projects can further customize the feature to provide users default issue templates for common communication such as bug reporting or feature requests [^3] https://docs.github.com/en/communities/using-templates-to-encourage-useful-issues-and-pull-requests/configuring-issue-templates-for-your-repository. Together, the name, advertised purpose, and user defaults normatively discourage appreciation. With hesitation, motivated users still reappropriate GitHub’s features to convey appreciation, as visible in issue #2264 in the react boilerplate project [^4] https://github.com/react-boilerplate/react-boilerplate/issues/2264, which is titled: “I don’t know how to thank, and show my appreciation, to the contributors on a good way”. GitHub users have pointed out how the reaction palette available within issues and pull requests lacks emojis to convey the sentiment of “Thank you” [^5] https://github.com/orgs/community/discussions/38201. While stars can be used to convey appreciation, they are also intended to function as a bookmarking tool and a way for the platform to learn user preferences [^6] https://docs.github.com/en/get-started/exploring-projects-on-github/saving-repositories-with-stars, which obscures the intention behind ‘starring’.
@@ 2.3. Appreciation systems in peer production
Our approach departs from prior attempts to support appreciation in open source by explicitly accounting for the manner in which software is ultimately used in the development environment. Prior work has studied several systems through which users of open source software can convey appreciation (Overney et al., 2020; Zhang et al., 2022b). This includes donation platforms such as PayPal, Patreon, and OpenCollective which projects may link to from their repositories (Overney et al., 2020). Similarly, GitHub has the ‘Sponsors’(Shimada et al., 2022) feature through which users of the platform can sponsor individual developers. The ‘Say Thanks’ [^7] https://github.com/BlitzKraft/saythanks.io project provides a link that contributors can include to the repository of a project, and that users can visit to send messages of appreciation. However, much like the GitHub features we described in the previous sections, all of these appreciation systems are detached from where software use occurs—the development environment. As a result, they present many of the barriers that we described in the previous section. By developing an appreciation system that is fine-tuned to users’ development practices, we investigate the potential benefits of such an approach as well as cross-cutting concerns that arise in its implementation. Developing an appreciation system that connects to the users’ development environment also alllows us to investigate how appreciation is experienced and expressed by users. This allows us to further extend prior literature because prior studies tend to focus solely on the experiences of the contributors receiving appreciation (Overney et al., 2020; Shimada et al., 2022)
@@
Our approach tries to enable users to express appreciation in terms of the abstractions they are exposed to—the package or its modules—rather than lower-level units of contribution such as individual commits or pull requests. This distinguishes it from some existing appreciation systems in peer production that restrict users to expressing appreciation towards individual units of work. Consider Wikipedia’s “Thanks” system (Matias et al., [n. d.]; Goel et al., 2019), through which editors can thank each other. A ”thank” link is shown next to each edit in the history view of an article. Clicking the link triggers a notification to the author of the edit. However, having to thank individual edits can be limiting when users want to express appreciation towards a group of edits, a paragraph, or a section. One comment on the talk page for Wikipedia’s “Thanks” system notes: “Sometimes I’d like to express thanks for a group of edits — for example, when none of them is individually a big deal, but together they’re really helpful. Any chance that we could get the chance to issue a single Thanks feature notice for a group of edits?” [^8] https://en.wikipedia.org/wiki/Help_talk:Notifications/Thanks/Archive_2. Our work adapts what users can thank to the form in which they consume the artifact (packages/modules) rather than the unit of production (pull requests or individual commits). We investigate the merits of such adaptation and reveal the constraints and requirements it entails. This allows us to derive implications for designing appreciation systems in other peer production contexts where units of work may be misaligned with what users feel appreciation towards.